erleliS 5s fe in tl estvation afact o suitch es are clases wken 1
2
erleliS 5s fe in tl estvation afact o suitch es are clases wken 1
2
Answer #1

When switch S1 is closed

C1 and C3 are in series

1/C13 = 1/1 + 1/3

C13 = 3/4 uF

C2 and C4 are in series

1/C24 = 1/2 + 1/4

C24 = 4/3 uF

and

above two combinations are in parallel

Ceq = C13 + C24

Ceq = 3/4 + 4/3 = 25/12 uF

so,

total charge

Q = Ceq * V

Q = 25 / 12 * 12

Q = 25 uC

Therefore,

Q (C1 ) = 9 uC

Q ( C2 ) = 16 uC

Q (C3 ) = 9 uC

Q (C4) = 16 uC

----------------------------------------------------------------------------------------------------

Now When both switches are closed

C1 and C2 are in parallel

C3 and C4 are in parallel

so,

C12 = 3 uF

C34 = 7 uF

This is in series

so,

1/Ceq = 1/3 + 1/7

Ceq = 21/10 uF

so,

total charge

Q = 21 / 10 * 12

Q = 25.2 uC

Potential across C1 and C2 = Q/C = 25.2 / 3 = 8.4 V

Potential across C3 and C4 = 25.2 / 7 = 3.6 V

so,

Q (C1 ) = 8.4 uC

Q ( C2 ) = 16.8 uC

Q (C3 ) = 10.8 uC

Q (C4) = 14.4 uC

------------------------------------------------------------------------------------------

In part (b), what is pent up ???? bad handwriting ....i am not able to read....However, I think you will be able to solve this with the above solution.

Please let me know if you have any doubt.